### v2

@@ A.2 EMCS Method for Chemical Kinetics
Given a specific reactor and different initial temperatures, pressures, species concentrations, the trajectories of all species form a low-dimensional manifold [45]. Manifold sampling comprises two parts. First, we sample from zero-dimensional homogeneous reactions, where heated, homogeneous fuel and oxygen ignite spontaneously. The initial temperature ($T$) range is set so that the ignition delay time belongs to $[0.1{\rm ms},10{\rm ms}]$. Initial pressure $P$ belongs to $[0.5{\rm atm},2{\rm atm}]$. Equivalence ratio $\phi$ belongs to $[0.5,3]$. We use the open-source chemical kinetics library Cantera [^1] https://cantera.org/ to run approximately 5000 cases with random initial conditions and maximal time step size $10^{-8}~{}{\rm s}$ until maximal running time (e.g., $10{\rm ms}$) or when the temperature change is smaller than 0.001 within $10^{-7}~{}{\rm s}$. We sample each state $\boldsymbol{x}(t)=[T(t),P(t),Y_{1}(t),\cdots,Y_{n_{s}}(t)]$ ($Y_{i}\in[0,1]$ is the mass fraction of species $i$) on each trajectory every $10^{\scriptscriptstyle{-7}}{\rm s}$ and its corresponding label $\boldsymbol{u}(t)=\boldsymbol{x}(t+\Delta t)-\boldsymbol{x}(t)$, where $\Delta t=10^{-6}{\rm s}$ is the step size we use for DNN model. The other is sampling from one-dimensional freely propagating premixed flame example at 1 atm, $\phi=1$, and temperature is set from 300K to 1000K (every 10K). The physical length is 0.03 m with 800 grid points. We sample from the stable states of grids. Both zero-dimensional and one-dimensional data constitute the manifold sample set, which is denoted as $\mathbb{D}_{\rm Init}=\{(\boldsymbol{x}(t),\boldsymbol{u}(t))\}$.
@@ A.3 Computational Singular Perturbation
and $f_{M+1}$ associated with $\tau_{M+1}$ mentioned in Eq. 13 is the fasted one within slow modes. To decouple fast modes and slow modes, we need to identify $\tau_{M+1}$. The motivation is that the integration of $\boldsymbol{\omega}_{\text{fast}}$ over time scale $\tau_{M+1}$ is negligible, namely,
@@ B.1 Details about Electronic Dynamical Process
In the EMCS stage, we first define the initial condition range as $U_{i},I_{i}\in[-10^{-3},10^{-3}]$ for manifold sampling, and randomly select 800 initial state vectors. Each state is then simulated for $1~{}\rm{ms}$, with 1,000 data points collected at a time interval of $10^{-6}~{}\rm{s}$. All the data points are combined and referred to as $\mathbb{D}_{\text{MF}}$. For the MC sampling, the range is set as $I_{1}\in[-10^{-5},10^{-7}]$, $I_{2}\in[-10^{-6},10^{-6}]$, with the remaining components in $[-10^{-3},10^{-3}]$. Subsequently, 5,000 points are randomly sampled and simulated for $k=30$ steps. Each evolution time $\tau_{j},0\leq j\leq k$, is randomly selected from $\tau_{j}\in[0,0.02]~{}\rm{s}$. Finally, all data are filtered based on the range of temporal change $[\lambda_{1}u^{\min}_{i},\lambda_{2}u^{\max}_{i}]$ over $\mathbb{D}_{\text{MF}}$, where $\lambda_{1}=\lambda_{2}=10$. Ultimately, we obtain the training dataset $\mathbb{D}_{\text{EMCS}}$.
We further present the comparison between DNN-MC/DeePODE predictions and direct integration results for $U_{2}$ and $I_{2}$. Our findings reveal that, in the high-frequency oscillation range of the ring modulator, DNN-MC faces difficulties in making accurate predictions, whereas DeePODE consistently provides accurate results. This demonstrates that DeePODE effectively captures the multiscale dynamics inherent in the system.
@@ B.2.1 Governing Equation
In Eq. [23], the transport of energy is formulated in terms of the total sensible enthalpy $h_{s}+\frac{1}{2}\vec{u}\cdot\vec{u}$, $h_{s}$ is the sensible enthalpy of mixture for ideal gases, $h_{s}=\sum_{k=1}^{N}Y_{k}h_{s,k}$, where $h_{s,k}=h_{k}-\mathring{h}_{k}$, $\mathring{h}_{k}=h_{k}(298{\rm K})$ is the enthalpy of formation of the $k$-th species. The divergence of the negative energy flux $-\nabla\cdot\dot{\vec{q}}$:
$-\nabla\cdot\dot{\vec{q}}=\nabla\cdot(\frac{\lambda}{c_{p}}\nabla h_{s})-\sum_{k=1}^{N}\nabla\cdot(\frac{\lambda}{c_{p}}h_{s,k}\nabla Y_{k})-\sum_{k=1}^{N}\nabla\cdot(h_{s,k}\hat{\vec{j}}_{k}),$ (28)
@@
$-\nabla\cdot\dot{\vec{q}}=\nabla\cdot(\lambda\nabla T)-\sum_{k=1}^{N}\nabla\cdot(h_{s,k}\hat{\vec{j}}_{k}).$ (30)
@@ B.2.2 Chemical Mechanisms
GRI-Mech 3.0, created by the Berkley combustion team, is a detailed kinetic mechanism for methane/air combustion including nitric species for NOx predictions with 53 species and 325 reactions  [^2] Gregory P. Smith, David M. Golden, Michael Frenklach, Nigel W. Moriarty, Boris Eiteneer, Mikhail Goldenberg, C. Thomas Bowman, Ronald K. Hanson, Soonho Song, William C. Gardiner, Jr., Vitali V. Lissianski, and Zhiwei Qin, http://combustion.berkeley.edu/gri-mech/version30/text30.html.. It is challenging to obtain a surrogate model of GRI-Mech 3.0 because of the large number of species. For example, [47] used an on-the-fly scheme to overcome this difficulty but with large computation cost, i.e., frequently retraining the neural network with data from the current simulation.
@@ C.1 Solver Indicator
where $\phi$ denotes the probability density function of the standard normal distribution, and $\mu(X)$,$\sigma(X)$ is the mean and standard deviation, respectively. The smaller is $p(y,X)$, the further distance is data from the training set. The confidence interval is dependent on the DNN model.
@@ C.2 Additional Characteristic Time Analysis
For the ring modulator model, the range for the MC method is defined as $I_{1}\in[-10^{-5},10^{-7}]$, $I_{2}\in[-10^{-6},10^{-6}]$, with the remaining components in $[-10^{-3},10^{-3}]$. For convenience, we set the EMCS evolution time sequence as $\boldsymbol{\tau}=[10^{-13},10^{-12},10^{-10},10^{-9},10^{-9},10^{-8},10^{-8},10^{-7},10^{-7}]~{}\rm{s}$. The hyper-parameters for the CSP method are set as $\text{tol}_{\text{rel}}=10^{-3}$ and $\text{tol}_{\text{abs}}=10^{-5}$. We then present the characteristic timescale distribution for each EMCS step. In Fig. 13(B), it can be observed that the EMCS method effectively covers a broad range of timescales.

### v3

@@ A.2 EMCS Method for Chemical Kinetics
Given a specific reactor and different initial temperatures, pressures, species concentrations, the trajectories of all species form a low-dimensional manifold [46]. Manifold sampling comprises two parts. First, we sample from zero-dimensional homogeneous reactions, where heated, homogeneous fuel and oxygen ignite spontaneously. The initial temperature ($T$) range is set so that the ignition delay time belongs to $0.1{\rm ms},10{\rm ms}]$. Initial pressure $P$ belongs to $0.5{\rm atm},2{\rm atm}]$. Equivalence ratio $\phi$ belongs to $[0.5,3]$. We use the open-source chemical kinetics library Cantera [^3] https://cantera.org/ to run approximately 5000 cases with random initial conditions and maximal time step size $10^{-8}~{\rm s}$ until maximal running time (e.g., $10{\rm ms}$) or when the temperature change is smaller than 0.001 within $10^{-7}~{\rm s}$. We sample each state $\boldsymbol{x}(t)=[T(t),P(t),Y_{1}(t),\cdots,Y_{n_{s}}(t)]$ ($Y_{i}\in[0,1]$ is the mass fraction of species $i$) on each trajectory every $10^{\scriptscriptstyle{-7}}{\rm s}$ and its corresponding label $\boldsymbol{u}(t)=\boldsymbol{x}(t+\Delta t)-\boldsymbol{x}(t)$, where $\Delta t=10^{-6}{\rm s}$ is the step size we use for DNN model. The other is sampling from one-dimensional freely propagating premixed flame example at 1 atm, $\phi=1$, and temperature is set from 300K to 1000K (every 10K). The physical length is 0.03 m with 800 grid points. We sample from the stable states of grids. Both zero-dimensional and one-dimensional data constitute the manifold sample set, which is denoted as $\mathbb{D}_{\rm Init}=\{(\boldsymbol{x}(t),\boldsymbol{u}(t))\}$.
@@ A.3 Computational Singular Perturbation
and $f_{M+1}$ associated with $\tau_{M+1}$ mentioned in Eq. 13 is the fastest one within slow modes. To decouple fast modes and slow modes, we need to identify $\tau_{M+1}$. The motivation is that the integration of $\boldsymbol{\omega}_{\text{fast}}$ over time scale $\tau_{M+1}$ is negligible, namely,
@@ B.1 Details about Electronic Dynamical Process
In the EMCS stage, we first define the initial condition range as $U_{i},I_{i}\in[-10^{-3},10^{-3}]$ for manifold sampling, and randomly select 800 initial state vectors. Each state is then simulated for $1~\rm{ms}$, with 1,000 data points collected at a time interval of $10^{-6}~\rm{s}$. All the data points are combined and referred to as $\mathbb{D}_{\text{MF}}$. For the MC sampling, the range is set as $I_{1}\in[-10^{-5},10^{-7}]$, $I_{2}\in[-10^{-6},10^{-6}]$, with the remaining components in $10^{-3},10^{-3}]$. Subsequently, 5,000 points are randomly sampled and simulated for $k=30$ steps. Each evolution time $\tau_{j},0\leq j\leq k$, is randomly selected from $\tau_{j}\in[0,0.02]~\rm{s}$. Finally, all data are filtered based on the range of temporal change $\lambda_{1}u^{\min}_{i},\lambda_{2}u^{\max}_{i}]$ over $\mathbb{D}_{\text{MF}}$, where $\lambda_{1}=\lambda_{2}=10$. Ultimately, we obtain the training dataset $\mathbb{D}_{\text{EMCS}}$.
We further present the comparison between DNN-MC/DeePODE predictions and direct integration results for $U_{2}$ and $I_{2}$ (Fig.  7). Our findings reveal that, in the high-frequency oscillation range of the ring modulator, DNN-MC faces difficulties in making accurate predictions, whereas DeePODE consistently provides accurate results. This demonstrates that DeePODE effectively captures the multiscale dynamics inherent in the system.
@@ B.2.1 Governing Equation
In Eq. [23], the transport of energy is formulated in terms of the total sensible enthalpy $h_{s}+\frac{1}{2}\vec{u}\cdot\vec{u}$, $h_{s}$ is the sensible enthalpy of mixture for ideal gases, $h_{s}=\sum_{k=1}^{N}Y_{k}h_{s,k}$, where $h_{s,k}=h_{k}-\mathring{h}_{k}$, $\mathring{h}_{k}=h_{k}(298{\rm K})$ is the enthalpy of formation of the $k$-th species. The divergence of the negative energy flux $\nabla\cdot\dot{\vec{q}}$:
$\nabla\cdot\dot{\vec{q}}=\nabla\cdot(\frac{\lambda}{c_{p}}\nabla h_{s})-\sum_{k=1}^{N}\nabla\cdot(\frac{\lambda}{c_{p}}h_{s,k}\nabla Y_{k})-\sum_{k=1}^{N}\nabla\cdot(h_{s,k}\hat{\vec{j}}_{k}),$ (28)
@@
$\nabla\cdot\dot{\vec{q}}=\nabla\cdot(\lambda\nabla T)-\sum_{k=1}^{N}\nabla\cdot(h_{s,k}\hat{\vec{j}}_{k}).$ (30)
@@ B.2.2 Chemical Mechanisms
GRI-Mech 3.0, created by the Berkeley Combustion Team, is a detailed kinetic mechanism for methane/air combustion including nitric species for NOx predictions with 53 species and 325 reactions  [^4] Gregory P. Smith, David M. Golden, Michael Frenklach, Nigel W. Moriarty, Boris Eiteneer, Mikhail Goldenberg, C. Thomas Bowman, Ronald K. Hanson, Soonho Song, William C. Gardiner, Jr., Vitali V. Lissianski, and Zhiwei Qin, http://combustion.berkeley.edu/gri-mech/version30/text30.html.. It is challenging to obtain a surrogate model of GRI-Mech 3.0 because of the large number of species. For example, [48] used an on-the-fly scheme to overcome this difficulty but with large computation cost, i.e., frequently retraining the neural network with data from the current simulation.
@@ C.1 Solver Indicator
where $\phi$ denotes the probability density function of the standard normal distribution, and $\mu(X)$$\sigma(X)$ is the mean and standard deviation, respectively. The smaller is $p(y,X)$, the further distance is data from the training set. The confidence interval is dependent on the DNN model.
@@ C.2 Additional Characteristic Time Analysis
For the ring modulator model, the range for the MC method is defined as $I_{1}\in[-10^{-5},10^{-7}]$, $I_{2}\in[-10^{-6},10^{-6}]$, with the remaining components in $10^{-3},10^{-3}]$. For convenience, we set the EMCS evolution time sequence as $\boldsymbol{\tau}=[10^{-13},10^{-12},10^{-10},10^{-9},10^{-9},10^{-8},10^{-8},10^{-7},10^{-7}]~\rm{s}$. The hyper-parameters for the CSP method are set as $\text{tol}_{\text{rel}}=10^{-3}$ and $\text{tol}_{\text{abs}}=10^{-5}$. We then present the characteristic timescale distribution for each EMCS step. In Fig. 13(B), it can be observed that the EMCS method effectively covers a broad range of timescales.
@@
C.3 Range Estimation
To demonstrate that a limited number of trajectories can provide a reasonable estimation of the data range, we visualized the dynamics of the maximum and minimum values for each dimension relative to the number of sampled trajectories. As shown in Fig. 14, the boundaries of the estimated range converge rapidly with only a small number of samples. The results reveal that a small number of data points based on manifold sampling could effectively estimate the data range.
Figure 14: The estimated range of Step 1 in EMCS as the number of initial trajectories varies. The orange line represents the lower bound, and the blue line represents the upper bound.
C.4 Ablation Study for DRM19 System
To quantify the contributions of the two EMCS components—Monte Carlo (MC) sampling and evolution-augmented generation—we conducted ablation experiments using the DRM19 chemical kinetics system (Section 3.4). We evaluated four models: MC sampling alone, MC with evolution augmentation (MC + Evolution), Multi-scale Sampling (MS) [8], and MS with evolution augmentation (MS + Evolution), using the root mean square error (RMSE) of predicted temperatures as the metric. Fig. 15 (A) shows that baseline MC and MS models yield RMSEs of $\mathcal{O}(10^{0})$, while their evolution-augmented counterparts reduce RMSE to $\mathcal{O}(10^{-2})$, highlighting the critical role of evolution augmentation. Fig. 15 (B, C) further demonstrates that increasing evolution steps (e.g., from two to three) significantly improve both point-wise and long-term temporal prediction accuracy, with diminishing returns beyond a few steps. These results confirm that evolution augmentation is the dominant contributor to model accuracy, as it captures the system’s multiscale dynamics, while MC sampling ensures broad phase space coverage.
Figure 15: Ablation study of EMCS. To ensure fairness, all four models in subplot (A) use the same data size of $1.1\times 10^{7}$. The training set for the MC + Evolution model is generated by taking $10^{6}$ MC random samples and evolving them for 10 steps, while the training set for the MS + Evolution model is generated in a similar manner. The MC model in subplot (B) is consistent with that in subplot (A). The remaining models (labeled as model 1 through model 10) are generated by evolving $10^{6}$ MC initial samples for 1 to 10 steps, respectively.
C.5 Baseline Comparisons with Advanced Sampling Techniques
To validate the advantages of the EMCS framework, we compared it against three baseline sampling techniques using the DRM19 chemical kinetics system (Section 3.4): (1) direct sampling from 3D turbulent reactive flow simulations (3D-Sim) [50, 51], (2) sampling from 1D steady-state reaction-diffusion systems with random perturbations (1D-Pert)  [52, 53], and (3) Multi-scale Sampling (MS) [8]. Datasets were constructed as described in Section 3.4 and Appendix A.2, and DNNs were trained to predict temporal evolution under identical initial conditions (1 atm, 1600 K, stoichiometric equivalence ratio). Fig. 16 shows that the EMCS-trained model demonstrates unique accuracy in its predictions, significantly outperforming 3D-Sim and MS model and maintaining stability unlike 1D-Pert, which diverges after 0.5 ms. Fig. 17 illustrates that 3D-Sim, 1D-Pert, and 0D datasets form distinct sub-manifolds with limited overlap, restricting generalization. EMCS, however, provides compact, configuration-independent coverage of the phase space, enhancing robustness across 0D, 1D, 2D, and 3D scenarios, as validated in Section 3.4 (Figs. 3 and 4).
Figure 16: Temporal prediction of models trained with different baseline sampling methods.
Figure 17: The data distribution of different sampling methods.